40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 270 of 341  Not logged in ELOG logo
ID Date Author Type Category Subject
  3585   Fri Sep 17 17:35:35 2010 steveUpdatePEMcleaned up at 1Y1 PSL rack

Cleaned up cables on the top and bottom. Vacuumed both areas. We still have some remaining shading from the MOPA umbilical and more unknown BNC cables hanging around.

  3584   Fri Sep 17 14:55:01 2010 josephbUpdateCDSTook 5565 RFM card from IOVME to place in the new IOO chassis

I took the 5565 RFM card out of the IOVME machine to so I could put it in the new IO chassis that will be replacing it.  It is no longer on the RFM network.  This doesn't affect the slow channels associated with the auxilliary crate.

  3583   Fri Sep 17 12:11:42 2010 josephbUpdateCDSDowns update

In doing a re-inventory prior to the IOO chassis installation, I re-discovered we had a missing interface board that goes in an IO chassis.  This board connects the chassis to the computer and lets them talk to each other.  After going to Downs we remembered Alex had taken a possibly broken interface board back to downs for testing. 

Apparently the results of that testing was it was broken.  This was about 2.5 months ago and unfortunately it hadn't been sent back for repairs or a replacement ordered.  Its my fault for not following up on that sooner.

I asked Rolf what the plan for the broken one was.  His response was  they were planning on repairing it, and that he'd have it sent back for repairs today.  My guess the turn around time for that is on the order of 3-4 weeks (based on conversations with Gary), however it could be longer.  This will affect when the last IO chassis (LSC) can be made fully functional.  I did however pickup the 100 foot fiber cable for going between the LSC chassis and the LSC computer (which will be located in 1X3).

As a general piece of information, according to Gary the latest part number for these cards is OSS-SHB-ELB-x4/x8-2.0 and they cost 936 dollars (latest quote).

  3582   Fri Sep 17 03:32:11 2010 KojiUpdateSUSArrangement of the SUS towers

The day before yesterday, I was cleaning a flow bench in the clean room.

I found that one SOS was standing there. It is the SRM suspension.

I thought of the nice idea:

- The installed PRM is actually the SRM (SRMU04). It is 2nd best SRM but not so diiferent form the best one.
==> Use this as the final SRM

- The SRM tower at the clean room
==> Use this as the final PRM tower.
==> The mirror (SRMU03) will be stored in a cabinet.

- The two SOS towers will be baked soon
==> Use them for the ETMs

This reduces the unnecessary maneuver of the suspension towers.

  3581   Fri Sep 17 03:06:06 2010 KojiUpdateSUSSOS sent for baking

Two SOS suspensions for the ETMs were disassembled and packed for cleaning and baking by Bob.

These suspensions have been stored on the X end flow bench long years, and looked quite old.

They have some differences to the modern SOSs.

- The top suspension block is made of aluminum and had dog clamps to fix the wires.
- The side bars are not symmetric: the side OSEM can only be fixed at the right bar (left side in the picture).
- EQ stops were made of Viton.
- One of the tower bases seems to have finger prints (of Mike Zucker?).

I found that the OSEM plates had no play. We know that the arrangement of the OSEMs gets quite difficult
in this situation. Therefore the holes of the screws were drilled with the larger drill.

We decided to replace all of the screws to the new ones as all of the screws are Ag plated and got corroded
by silver sulfide (Ag2S). I checked our stock in the clean room. We have enough screws.

Important note: Use stainless screws in aluminum / Silver-plated screws in stainless
There exists some study about galling: LIGO-G020394-00-D

Attachment 1: IMG_3596.jpg
IMG_3596.jpg
Attachment 2: IMG_3597.jpg
IMG_3597.jpg
  3580   Fri Sep 17 01:36:14 2010 valeraUpdate PMC line width

The attached plots show the PMC cavity line width measurement with 1 mW and 160 mW into the PMC. The two curves on each plot are the PMC transmitted power and the ramp of the fast input of the NPRO. The two measurements are consistent within errors - a few %. The PMC line width  3.5 ms (FWHM) x 4 V / 20 ms (slope of the ramp) x 1.1 MHz / V (NPRO fast actuator calibration from Innolight spec sheet) = 0.77 MHz.

Here is the output of the calculation using Malik Rakhmanov code:

 

modematching =  8.4121e-01

transmission1 =   2.4341e-03

transmission2 =   2.4341e-03

transmission3 =   5.1280e-05

averageLosses =  6.1963e-04

visibility =  7.7439e-01

Here are the inputs for the calculation in the param.m:

 

fw = 0.77e6;                % width of resonance (FWHM) in Hz

Plas = 0.164;                % power into the PMC in W

 

% the following number refer to the in-lock cavity state

 

Pref = 0.037;                % reflected power in W

Ptr = 0.0712;                 % transmitted power in W

Pleak = 0.0015;              % power leaking from back of PMC in W

 

 

Attachment 1: TEK00009.PNG
TEK00009.PNG
Attachment 2: TEK00010.PNG
TEK00010.PNG
  3579   Wed Sep 15 19:29:13 2010 valeraSummary PSL power budget
 Location  Power (mW)
 NPRO - after HWP  252
 Rejected by input FI polarizer  38
 After output FI polarizer  175
 Into PMC  164
 PMC reflected  37
 PMC transmitted  71
 PMC leakage  1.5
 After PMC TRANS PD/Camera BS

 1.2

 After RefCav EOM  1.1
 Into RefCav  0.3

 Notes:

- NPRO injection current 1.0 A

- PMC losses ~32%

- FSS AOM diffraction efficiency ~52%

  3578   Wed Sep 15 16:12:35 2010 koji, steveUpdateMOPAMOPA Controller is taken out of the PSL rack

We removed the Lightwave MOPA Controller, PA#102, NPRO206 power supply to make room for the IOO chassy at 1X1 (south) rack.

The umbilical cord was a real pain to take out. It is shading its plastic cover. The unused Minco was disconnected and removed.

The ref. cavity ion pump controller- power supply was temporarily taken out also.

Attachment 1: P1060843.JPG
P1060843.JPG
  3577   Wed Sep 15 16:00:26 2010 koji, steveUpdateMOPA 

We removed the Lightwave MOPA Controller from 1X1 (south)  It was a real painfully messy job to pull out the umbilical.

Note: the umbilical is shading it plastic cover. It is functional but it has to be taken out side and cleaned. Do not remove it from it's plastic bag in a clean environment.

Now Joe has room for IOO chassy  in this rack.

We also removed the Minco temp controller and ref. cavity ion pump power supply.

 

  3576   Wed Sep 15 14:34:57 2010 josephbSummaryCDSPlan for RFM switch over

Steps for RFM switch over:

1) Ensure the new frame builder code is working properly:

   A) Get Alex to finish compiling the frame builder and test on Megatron.

   B) Test the new frame builder code on fb40m (which is running Solaris) in a reversible way.  Change directory structure away from Data1, Data2, to use actual times.

   C) Confirm new frame builder code still records slow channels (c1dcuepics).

2) Ensure awg, tpman, and diagnostic codes (dtt) are working with the new front end code.

3) Physically move RFM cables from old front ends to the new front ends.  Remove excess connections from the network.

4) Merge the megatron/c1sus/c1iscex/c1ioo network with the main network.

   A) Update all the network settings on the machines as well as Linux1

   B) Remove the network switch separating the networks.

4) Start the new frame builder code on fb40m.

  3575   Wed Sep 15 03:08:26 2010 KojiUpdatePSLFSS locking

Brilliant! This is the VERY way how the things are to be conquered!

Quote:

The RefCav is locked and aligned. I changed the fast gain sign by changing the jumper setting on the TTFSS board. The RefCav visibility is 70%. The FSS loop ugf is about 80 kHz (plot attached)  with FSS common gain max out at 30 dB. There is about 50 mW coming out of the laser and a few mW going to RefCav out of the back of the PMC. So the ugf can be made higher at full power. I have not made any changes to account for the PMC pole (the FSS is after the PMC now). The FSS fast gain was also maxed out at 30 dB to account for the factor of 5 smaller PZT actuation coefficient - it used to be 16 dB according to the (previous) snap shot. The RefCav TRANS PD and camera are aligned. I tuned up the phase of the error signal by putting cables in the LO and PD paths. The maximum response of the mixer output to the fast actuator sweep of the fringe was with about 2 feet of extra cable in the PD leg.

I am leaving the FSS unlocked for the night in case it will start oscillating as the phase margin is not good at this ugf.

 

  3574   Wed Sep 15 01:58:28 2010 valeraUpdatePSLFSS locking

The RefCav is locked and aligned. I changed the fast gain sign by changing the jumper setting on the TTFSS board. The RefCav visibility is 70%. The FSS loop ugf is about 80 kHz (plot attached. there is 10 dB gain in the test point path. this is why the ugf is at 10 dB when measured using in1 and in2 spigots on the front of the board.)  with FSS common gain max out at 30 dB. There is about 250 mW coming out of the laser and 1 mW going to RefCav out of the back of the PMC. So the ugf can be made higher at full power. I have not made any changes to account for the PMC pole (the FSS is after the PMC now). The FSS fast gain was also maxed out at 30 dB to account for the factor of 5 smaller PZT actuation coefficient - it used to be 16 dB according to the (previous) snap shot. The RefCav TRANS PD and camera are aligned. I tuned up the phase of the error signal by putting cables in the LO and PD paths. The maximum response of the mixer output to the fast actuator sweep of the fringe was with about 2 feet of extra cable in the PD leg.

I am leaving the FSS unlocked for the night in case it will start oscillating as the phase margin is not good at this ugf.

Attachment 1: DSC_2510.JPG
DSC_2510.JPG
  3573   Wed Sep 15 01:27:52 2010 rana, steve, valeraUpdatePSLFSS cables connected

- connected the TTFSS cables (FSS fast goes directly to NPRO PZT for now)

- measured the reference cavity 21.5 MHz EOM drive to be 17.8 dBm

-  turned on the HV for the FSS phase correcting EOM (aka PC) drive

- connected and turned on the reference cavity temperature stabilization

- connected the RefCav TRANS PD

- fine tuned the RefCav REFL PD angle

  3572   Tue Sep 14 18:07:41 2010 AlbertoUpdateElectronicsFrequency Box Documentation Added to the SVN

I completed a LIGO document describing design, construction and characterization of the RF System for the 40m upgrade.

It is available on the SVN under https://nodus.ligo.caltech.edu:30889/svn/trunk/docs/upgrade08/RFsystem/RFsystemDocument/

It can also be found on the 40m wiki (http://lhocds.ligo-wa.caltech.edu:8000/40m/Upgrade_09/RF_System#preview), and DCC under the number T1000461.

  3571   Tue Sep 14 00:21:51 2010 ranaOmnistructureEnvironmentOffice area temperature change

I changed the setpoint for the HVAC control (next to Steve) from 73F to 72F. This is to handle the temperature increase in the control room with the AC unit there turned off.

We know that the control setpoint is not linear, but I hope that it settles down after several hours. Lets wait until Tuesday evening before making another change.

  3570   Mon Sep 13 22:51:07 2010 tara,valeraConfigurationPSLbeam scan for RCAV

On Friday, Valera and I calculated the modematching for reference cavity from AOM.

We scan the beam profile where the spot should be.

The first beam waist in the AOM is 103 um, the lens (f= 183 mm, I'm not sure if I have the focal length right) is 280 mm away.

The data is attached. The first column is marking on the rail in inches,

the second column is distance from the lens, the third and fourth column are

vertical and horizontal spot radius in micron. Note that the beam is very elliptic because of the AOM.

Attachment 1: 2010_09_10_w.mat
  3569   Mon Sep 13 21:52:53 2010 kiwamuUpdateGreen Lockingthe X end laser is ON

 I turned ON the laser at the X end station, which had been OFF for several weeks because of the crane business.

Now the green beam hits the ITMX and I got a reflection back to the end table.   

This green beam will be a nice reference when we install the green periscope in the chamber.

If it's necessary, feel free to correct the alignment of the green beam during my absence.

  3568   Mon Sep 13 19:41:38 2010 ranaUpdatePSLFSS AOM alignment

The IR sensitive Olympus 570 camera gives us a really nice view of these IR beams. Its actually a lot better than what you can get with the analog IR viewers:

 

PSLAOMdogs
  3567   Mon Sep 13 12:09:42 2010 kiwamuSummaryGeneralelog restarted

 Elog didn't respond, so I restarted it with the script.

Before killing the process somehow two elog daemons were running at the same time. I killed those two manually. 

  3566   Mon Sep 13 11:49:34 2010 kiwamuUpdatePSLmode matching from PMC to IMC

I have been working on the mode matching lenses which are sitting after the boradband EOM.

Last Friday I checked the mode profile after the first mode matching lens (f=-150mm). The measured mode was good.

According to the calculation done by ABCD software, the waist size is supposed to be 80.9 um after that lens.

The measured waists are 80.5 um for the vertical mode and 79.4 um for the horizontal mode.

The screenshot of the ABCD's result and the plot for the mode measurement are shown below.

I didn't  carefully check the mode after the last convex lens (f=200mm), but it must be already good because the beam looks having a long rayleigh range.

Now the beam is reflected back from MC1 and goes to the AP table since I coarsely aligned the beam axis to the MC.

newMM.png  psl_mmt1.png

 

/****  fitting result ****/

w0_v =  80.4615      +/- 0.1695 [um] 

w0_h =  79.4468      +/- 0.1889 [um]

z_v =  -0.115234        +/- 0.0005206  [m]

z_h =  -0.109722        +/- 0.0005753 [m]

 

  3565   Mon Sep 13 11:40:50 2010 AlbertoUpdateElectronicsFrequency Box Documentation Added to the SVN

I uploaded all the material about the RF frequency Generation Box into the SVN under the path:

https://nodus.ligo.caltech.edu:30889/svn/trunk/docs/upgrade08/RFsystem/frequencyGenerationBox/

I structured the directory as shown in this tree:

freqBoxSVNdierctoryStructure.png

I'm quickly describing in a section of the Rf system upgrade document with LIGO # T1000461.

  3564   Mon Sep 13 10:22:48 2010 josephbUpdateCDSRCG bugs/feature request wiki page

I've started a wiki page under the Upgrade 09/New CDS section regarding known bugs and pending feature requests for the Real Time Code Generator.   It can be found at http://lhocds.ligo-wa.caltech.edu:8000/40m/Bugs_and_Pending_Feature_requests_for_the_RCG.  If you have any ideas to improve the RCG or encounter a bug in the code generation process (say a particular part doesn't work inside subsystems for example), please note them here.

Currently there are bugs with excitation points (they don't work inside of a subsystem block) and tags (they don't respect scope and only 1 "from" tag for each "goto" tag connected to the output of a subsystem block).

  3563   Mon Sep 13 02:45:59 2010 ranaConfigurationDMFseisBLRMS restarts

I restarted the seisBLRMS DMF monitor by ssh'ing into mafalda and starting up a matlab session. I also have started a StripTool session on rossa by forwarding the process from op440m.

We need to get the modern EPICS installation onto these linux machines by copying what K. Thorne has done at LLO.

  3562   Mon Sep 13 00:19:32 2010 ranaUpdate VCO Driver Output power v. slider control voltage

I measured the RF power output of the VCO Driver box as a function of slider value. I measured using the Gigatronics Handheld power meter and connected to the AOM side of the cable after the white Pasternak DC block.

* at low power levels, I believe the waveform is too crappy to get an accurate reading - that's probably why it looks non-monotonic.

* the meter has a sticker label on it saying 'max +20 dBm'. I went above +20 dBm, but I wonder if maybe the thing isn't linear up there...

Attachment 1: vco.png
vco.png
  3561   Sun Sep 12 23:16:52 2010 valeraUpdate FSS mode matching

The attached plot shows the beam scans of the beam leaking from the back mirror of the PMC to the BS cube that first turns the S-pol beam 90 deg to the AOM and then transmits the AOM double passed and polarization rotated P-pol beam to the reference cavity. The beam from the PMC is mode matched to the AOM using a single lens f=229 mm. The ABCD file is attached. The data were taken with VCO control voltage at 5 V. We then reduced the voltage to 4 V to reduce the astigmatism. Tara has the data for the beam scan in this configuration in his notebook.

The beam from AOM is mode matched to the reference cavity using a single lens f=286.5 mm. The ABCD file is attached.

Attachment 1: fss.pdf
fss.pdf
Attachment 2: fssaom-abcd.tiff
Attachment 3: fssrc-abcd.tiff
  3560   Sun Sep 12 23:02:53 2010 valeraUpdate PMC mode matching

Kiwamu and I found that the first lens in the PMC mode matching telescope was mislabeled. It is supposed to be PLCX-25.4-77.3-C and was labeled as such but in fact it was PLCX-25.4-103.0-C. This is why the PMC mode matching was bad. We swapped the lens for the correct one and got the PMC visibility of 82%. The attached plot shows the beam scans before and after the PMC. The data were taken with the wrong lens. The ABCD model shown in the plot uses the lens that was there at the time - PLCX-25.4-103.0-C. The model for the PMC is just the waist of 0.371 mm at the nominal location. The snap shot of the ABCD file is attached. The calculation includes the KTP for FI and LiNb for EOM with 4 cm length. The distances are as measured on the table.

Attachment 1: pmc.pdf
pmc.pdf
Attachment 2: pmc-abcd.tiff
  3559   Sun Sep 12 22:36:03 2010 ranaConfigurationComputersrossa notes

I found Sanjit's instructions for doing the Nvidia settings too complicated and so I followed these instructions from Facebook:

http://www.facebook.com/notes/centos-howtos/installing-nvidia-display-drivers-on-centos-55/399295987425

After installations, the monitors were autodetected and the Xinerama effect is working.

  3558   Sat Sep 11 22:42:07 2010 valeraUpdatePSLPSL update

- The PMC REFL PD was moved from the temporary location to the one called for by the PSL layout (picture attached). The leakage beams were dumped.

- The FSS reference cavity was aligned using temporary periscope and scanned using NPRO temperature sweep. The amplitude of the sweep (sine wave 0.03 Hz) was set such that the PMC control voltage was going about 100 V p-p with. With rough alignment the visibility was as high as 50% - it will be better when the cavity is locked and better aligned but not better than 80% expected from the mode astigmatism that Tara and I measured on Thursday. The astigmatism appear to come from the FSS AOM as it depends on the AOM drive. We reduced the drive control voltage from 5 V to 4V beyond that the diffraction efficiency went below 50%. The FSS REFL PD was set up for this measurement as shown in the attached picture. There is also a camera in transmission not shown in the picture.

Attachment 1: DSC_2502.JPG
DSC_2502.JPG
Attachment 2: DSC_2505.JPG
DSC_2505.JPG
  3557   Sat Sep 11 03:16:51 2010 ranaConfigurationComputersrossa notes

I wiped out the old CentOS install on rossa and installed CentOS 5.5 on there. The DVDs are on a spindle in the control room; there were 2 iso's, but I only needed the first to install most things.

It still needs to get all of the usual stuff (java, flash, nvidia) installed as well as setting up the .cshrc and the NFS mount of /cvs/cds. The userID and groupID are set to 1001 as before. Whoever

sees Sanjit first should steer him towards this elog entry.

 

  3556   Fri Sep 10 02:35:30 2010 rana, kiwamu, valera, taraConfigurationPSLupdate

Over the last couple nights we got the beam into the FSS path and all the way to the IMC and out onto the AP table.

Tara and Valera have calculated a mode-matching solution for the reference cavity. It utilizes only a single lens between the AOM and the reference cavity. Valera and Steve will move the reference cavity into place in the morning.

We noticed that the layout was too tight on the end going into the MC and so we adjusted the angle of the final zig-zag. This will put the final mode-matching lens in between the final steering mirrors (which is generally undesirable) but the lens in this case is only f=400 mm. In addition, this lens may provide some more decoupling between the steering mirrors.

The whole layout has to be a little adjusted because of a calculation mistake I made in the mode-matching. I used only the nominal focal lengths from the CVI catalog and not the effective one. For the UV-grade fused silica lenses, the effective focal length is actually 20-30% longer. Today we measured that the "f=200 mm" lens we got is actually f = 238 mm. The BK7 lenses are much closer to the nominal.

We also replaced the Klinger mount ahead of the PMC with a Polanski style so that we could get the PMC REFL beam out without hitting the mount. Valera will continue to refine this section on the weekend.

Tomorrow, we will lock the MC using feedback only to the NPRO. The 0-150 V piezo driver is on the PSL table ready for action.

I also got a LCD video monitor from Frank and hooked it up on the PSL table. If we like this kind of thing, we can get many of them. They are pretty cheap. It would be handy to have 3-4 of them on the PSL and one on every of the ISC tables. They take the standard video for input and need +12V for power. Right now the one in there is looking at the PMC transmission.

The Omnigraffle layout as of tonight is attached.

Attachment 1: 40mPSL.pdf
40mPSL.pdf
  3555   Thu Sep 9 18:53:56 2010 AlbertoConfigurationElectronicsBusby Box, Rai's Box, SR554 in the RF cabinet

I stored the Busby Box, the Rai's Box and the SR554 preamp in the RF cabinet down the Y arm.

  3554   Thu Sep 9 16:54:56 2010 ranaUpdateSAFETYmain laser is back ON

I turned off the laser last night to protect against the electricians. but failed to elog it. I accept the burnt toast completely.

  3553   Thu Sep 9 14:42:40 2010 josephbUpdateCDSUpdating suspensions model

I've been working on updating the suspensions model, incorporating Koji's refinements as well as trying to simplify the model and making it less cluttered. 

This had the side benefit of making an incorrect connection obvious.  I had incorrectly wired the ASCPIT input to be summed into the yaw path, and the ASCYAW input into the pitch path.  This has been corrected.

I've finished a single optic, and now I am in the process of propagating the changes to all the optics, as well as cleaning up the overall diagram using Rolf's new tags, which make things much less cluttered.  I've attached a screenshot of the PRM optic control model, and will be updating the Matlab web export once I've updated the full model.

Attachment 1: SingleSUS.png
SingleSUS.png
  3552   Thu Sep 9 12:02:03 2010 AlbertoUpdateElectronicsFrequency Generation Box - Amplitude Noise Measurements

I measured the amplitude noise of the source outputs and the EOM outputs of the Frequency Generation box.

the setup I used is shown in this diagram:

FreqBoxAmplitudeNoiseMeasurementSetup.png

(NB It's important that the cables from the splitter to the RF and LO inputs of the mixers are the same length).

The results of the measurements are shown in the following plot:

 FreqBoxAmplitudeNoise_BusbyBox_AllChannels.png

Considerations:

1) both Crystals (29.5MHz and 11MHz) have the same noise

2) the 55MHz source's noise is bigger than the 11 MHz (~2x): the frequency multiplication and amplification that happen before it, add extra noise

3) the noise at EOM outputs is ~2x bigger than that of the relative sources

 

When I have the chance, I'll plot the results of my calculations of expected noise and compare them with the measurements.

  3551   Thu Sep 9 11:31:54 2010 steveUpdatePEMaccelerometer cable connection

I turned off/on the power to the accelerometers in order to re rout their connections. I found cable connector body-nut  #3 loose to Accelerometer 2X This connector should be checked for solid performance.

  3550   Thu Sep 9 11:17:15 2010 steveUpdateSAFETYmain laser is back ON

Quote:

Innolight 2W 1064nm main laser was turned off for more enclosure related work.

The electrician re rooting the power- conduit to the interlock under the floor and bring it through the new concrete tunnel.

Emergency laser power shut off switch was tested at entry door 104M.  It worked fine.

 

No laser glasses required. Laser will be turned back on around 11 am today.

Atm1,   AC power line to enclosure and interlock  was extended.

Atm2,   24V line to interlock rerouted

Atm3,   job is completed

 

SAFETY GLASSES REQUIRED !  SAFETY GLASSES REQUIRED!

 

 The laser is turned on. Injection current set to 0.871A.

Actually some one turned the  laser  off last night and did  not enter it into the elog ! Burned toast award is granted !

Please  come forward voluntarily.

Attachment 1: P1060813.JPG
P1060813.JPG
Attachment 2: P1060814.JPG
P1060814.JPG
Attachment 3: P1060816.JPG
P1060816.JPG
  3549   Thu Sep 9 08:23:45 2010 steveUpdateSAFETYmain laser turned off

Innolight 2W 1064nm main laser was turned off for more enclosure related work.

The electrician re rooting the power- conduit to the interlock under the floor and bring it through the new concrete tunnel.

Emergency laser power shut off switch was tested at entry door 104M.  It worked fine.

 

No laser glasses required. Laser will be turned back on around 11 am today.

  3548   Thu Sep 9 05:58:04 2010 kiwamuUpdatePSLmode matching from PMC to IMC

 I started mode matching of the beam going to IMC. The work is still going on.

According to Rana's calculation (see here), I put the first lens (f=200mm) in between two steering mirrors after PMC.

The distance from PMC to the first lens was adjusted by using a metal ruler.  So I believe the accuracy is something like 1mm.

I aligned the beam path going through the broadband EOM and the mode matching lenses.

 

  I could find the optimum position for the second lens  (f=-150mm) by sliding the position of the lens and measuring the mode after it.

But the optimum position looked a bit far from the EOM. It's off by about 3-4 inch from the designed position.

Somehow I feel that the beam before the second lens goes with a smaller divergence angle than that of designed.

So tomorrow I am going to restart the work from checking the mode before the second lens.

Maybe at first I should measure the mode without going through the EOM because it changes the waist position and makes the system not straightforward.

  3547   Wed Sep 8 16:17:58 2010 steveUpdatePEMparticle counter problem

I played with the particle counter this morning. Now it's running and reading the correct numbers on the old COCHECKLIST.adl medm screen.

However, it can not be seen with dataviewer.

Koji just rebooted the frame builder and the problem was solved.

  3546   Wed Sep 8 14:44:37 2010 josephbUpdateCDSUpdated parse_mdl_to_ipc.py and feCodeGen.pl

I updated parse_mdl_to_ipc.py to correctly work with the 3 new cdsIPCx parts, namely cdsIPCx_PCIE (for dolphin connections), cdsIPCx_RFM (for our traditional reflected memory connections), and cdsIPCx_SHMEM (local shared memory on the computer).  These parts replaced cdsIPCx awhile back (see here ).

The code now correctly counts each type independantly with regards to ipcNum (I.e. you can have ipcNum = 0 for RFM and ipcNum = 0 for SHMEM for example).

 

I also went in and modified a few sections of the feCodeGen.pl (in /opt/rtcds/caltech/c1/core/advLigoRTS/src/epics/util/) so as to properly assign names to matrix adl files, matrix of filter bank adl files, and filter bank adl files.

  3545   Wed Sep 8 11:56:24 2010 kiwamuSummaryCDSSeptember CDS test plan

 Joe and Kiwamu

We discussed about our CDS plan for this September. The summary of the plan and "to do list" are now on the wiki page;

http://lhocds.ligo-wa.caltech.edu:8000/40m/Upgrade_09/CDS/September_CDS_plan

 

  Basically there are three major missions that we will do in this month;

1. complete damping of the vertex suspensions
2. Preparation for Green locking
3. Development of Simulated Plants
 

 We also try to keep updating the wiki page.

  3544   Wed Sep 8 11:46:53 2010 kiwamuUpdatePSLupdate of the layout

I put some green stuff on the layout drawing.

I continue to refine the positions of these stuff.

 

Notes :

 1. I flipped the reference cavity. So now the cavity is sitting on the left hand side of the layout.

  2. I removed the ISS stuf. We should think about where ISS should be.

Quote:

Also, Kiwamu has modified the layout drawing to add the green PLL stuff. This has collapsed the reference cavity's wave function placing it close to its original position.

WE (maybe Valera and Steve) can now put the reference cavity back on the table.

 

Attachment 1: 40mPSL.png
40mPSL.png
  3543   Wed Sep 8 09:50:13 2010 ranaUpdatePSLPSL update

Also, Kiwamu has modified the layout drawing to add the green PLL stuff. This has collapsed the reference cavity's wave function placing it close to its original position.

WE (maybe Valera and Steve) can now put the reference cavity back on the table.

  3542   Wed Sep 8 00:01:07 2010 rana, valeraUpdatePSLPMC update

We ran the cables for the PMC: The RF cable for the 35.5 MHz drive was cheap and so we swapped the 29.5 MHz cable for it.

There now remain 1 RG-174 cable to drive the FSS PC (21.5 MHz) and 3 Heliax for the Kiwamu Tri-Mod EOM (11, 29.5, and 55 MHz).

We also changed the BLACK HV drive cable for the RED one (previously used for the MZ). All HV cables MUST be RED.

The BLACK cable is now used for the PMC_REFL DC.

The Heliax cables are routed onto the table - it remains a Alberto/Kiwamu job to strain relieve them and attach them to the TriMod box and EOM in the morning.

The PMC is locked and we did some partially bootless alignment and mode-matching. It locks easily on a TEM00 mode (with very poor visibility), but the

rest of the beam train can now be aligned while Valera does the PMC matching mambo.   

 

  3541   Tue Sep 7 23:49:08 2010 sanjitConfigurationComputersaldabella network configuration

 

added name server 192.169.113.20 as the first entry in /etc/resolv.conf

changed the host IPs in /etc/hosts to 192.168.xxx.yyy

made:

127.0.0.1 localhost.localdomain localhost

::1 localhost6.localdomain6 localhos6

as the first two lines of /etc/hosts

 

/cvs/cds mounts

on ethernet, DNS look-up works without the explicit host definitions in /etc/hosts,

but those entries are needed for wifi only connection.

 

  3540   Tue Sep 7 23:34:15 2010 Kiwamu, SanjitConfigurationComputerse-log

e-log was repeatedly hanging and several attempts to start the daemon failed.

problem was solved after clearing the (firefox) browser cache, cookie, everything!!

 

  3539   Tue Sep 7 23:17:45 2010 sanjitConfigurationComputersrossa notes

Quote:

* rossa needs to be able move windows between monitors: Xinerama?

 Xinerama support has been enabled on rossa using nvidia-settings.

  3538   Tue Sep 7 22:21:47 2010 DmassConfigurationelogelog restarted

Quote:

elog crashed on an upload. restarted and it worked fine with the same file.

 Again. Resubmitted an old entry with just text changes. Elog hung for 5 min +.

  3537   Tue Sep 7 22:21:17 2010 AlbertoConfigurationComputerselog restarted
  3536   Tue Sep 7 20:44:54 2010 YoichiHowToCOMSOL TipsCOMSOL example for calculating mechanical transfer functions

I added COMSOL example files to the 40m svn to demonstrate how to make transfer function measurements in COMSOL.

https://nodus.ligo.caltech.edu:30889/svn/trunk/comsol/MechanicalTF/

The directory also contains an (incomplete) explanation of the method in a PDF file.

ELOG V3.1.3-